# Composable Interventions for Language Models

### v1

@@ Abstract
Test-time interventions for language models can enhance factual accuracy, mitigate harmful outputs, and improve model efficiency without costly retraining. But despite a flood of new methods, different types of interventions are largely developing independently. In practice, multiple interventions must be applied sequentially to the same model, yet we lack standardized ways to study how interventions interact. We fill this gap by introducing composable interventions, a framework to study the effects of using multiple interventions on the same language models, featuring new metrics and a unified codebase. Using our framework, we conduct extensive experiments and compose popular methods from three emerging intervention categories—knowledge editing, model compression, and machine unlearning. Our results from 310 different compositions uncover meaningful interactions: compression hinders editing and unlearning, composing interventions hinges on their order of application, and popular general-purpose metrics are inadequate for assessing composability. Taken together, our findings showcase clear gaps in composability, suggesting a need for new multi-objective interventions.111All of our code is public: github.com/hartvigsen-group/composable-interventions††Correspondence: {arinbjorn,hartvigsen}@virginia.edu, kyobrien@microsoft.com
@@ 1 Introduction
Language models (LMs) exhibit striking capabilities on important tasks in myriad domains including medicine [1], finance [2], science [3], and entertainment [4]. But despite high performance, deployed LMs can still misbehave unpredictably and require updates. For example, LMs generate content that is hallucinatory [5], factually incorrect [6], and harmful [7, 8, 9]. Beyond unwanted behaviors, user requirements also change over time. For example, regulations arise [10], computational resources constrict, knowledge gets outdated [11], and copyrighted training materials are identified [12]. Without ways to quickly address these issues, models can be left miscalibrated, outdated, and biased, limiting their widespread responsible use [13].
@@
Many recent works study efficient, in-place updates for LMs. We broadly refer to these as interventions—updates to targeted properties of LMs applied after pretraining (and optional fine-tuning). For example, we can view model compression [14, 15, 16] as an intervention to make language models more inference- or memory-efficient. Other rapidly-advancing examples include knowledge editing [17, 18, 19, 20, 21, 22], detoxification [23, 24, 25, 26], and unlearning [27, 28, 29]. However, these interventions are largely advancing independently. In practice, we usually have multiple requirements for properties of our models (e.g., factuality and efficiency). And as requirements change over time, new interventions must be applied. While some works have started studying interactions between training objectives [30, 31, 32], practical widespread use is limited without unified evaluations for how interventions interact.
@@
As depicted in Figure 1, we propose that practical interventions should be composable: When an intervention is applied to a model, it should not interfere with prior or future interventions. For example, we might repair a factual error in an LM using a knowledge editing method. Later on, we might quantize that model. A composable quantization method should specifically preserve the edited fact and a composable editor should withstand quantization. However, it remains unknown how well existing interventions compose, and we lack formal notions of intervention composition. We therefore propose two metrics for composability: 1) Order-free Error, where an intervention is composable if its application leaves others’ success unimpacted, and 2) Order Sensitivity, where the combined success of multiple interventions should not depend on the order in which they are applied. We instantiate these metrics in a codebase that includes popular interventions, tackling a major lack of standardized code that has inhibited cross-intervention evaluations until now.
@@
We use our framework to extensively study composability of state-of-the-art knowledge editing, model compression, and machine unlearning interventions on Llama3-8B [33]. Our experiments with 310 different compositions unearth novel insights and offer guidance for composing recent methods. Three key results are as follows: Model compression often limits the success of other interventions. This suggests a significant drawback of existing compression methods, which are nearly universal in LM deployments. The order in which interventions are applied dramatically alters their success. Therefore, we need new interventions explicitly designed for composability. General-purpose post-intervention model performance is a poor proxy for composability. This indicates that targeting composability as a metric has the potential to drive the development of new, practically-grounded interventions. Taken alongside our other findings (Section 4), our experiments suggest we sorely need to broaden intervention method evaluations and to design new methods explicitly for composability.
@@
We introduce the notion of composability to language model interventions, formalizing the sequential applications of multiple interventions to LMs. Our work opens doors to address crucial, practical challenges in online LM updates and broadens evaluation criteria for interventions.
@@
2 An Evaluation Framework for Composable Interventions
@@ 2.1 Preliminaries: Interventions for Language Models
Assume we are given a language model , where contains its pre-trained parameters and is the parameter space. We define an intervention to be the combination of four things: 1) an operator that produces a new model , 2) hyperparameters , which typically control the strength of intervention, 3) a loss that specifies the targeted update,222Some interventions require data, some do not. We remove direct data dependencies for simplicity. and 4) evaluation criteria , where higher scores indicate better performance. Each intervention’s operator aims to minimize loss and ultimately achieve high performance according to . For readability, we denote the application of intervention to model as .
@@ 2.2 Composable Interventions for Language Models
Now consider possible interventions . We define intervention composition as the application of multiple interventions to a model in sequence.333For readability, we describe our framework in terms of two interventions. For example, we might apply intervention to model , then to the intervened-upon model. We can write this composition as either or , and here we use the latter format.
Most interventions are developed specifically for their own criterion . For simplicity, we let each intervention have one criterion, though in practice they often have multiple. We propose that the quality of many interventions in practice should also relate to how an intervention impacts our ability to apply other interventions. To meet this need, we propose two metrics that measure such composability. Each is computed for a given criterion and can be computed for a single hyperparameter setting or ranged over multiple settings to attain a richer description of intervention interactions.
@@
A smaller area again indicates better composability. And again, if hyperparameters are already chosen, this simplifies to a simple error: .
@@ 3.1 Intervention Datasets, Methods, and Metrics
We next detail the ten techniques selected from the three categories of interventions, along with their task-specific and overall utility metrics. Appendix B further describes the backgrounds of the three types of intervention and Table 1 summarizes our experimental setup.
| Interventions | Methods | Datasets | Intervention Metrics | Composability Metrics |
@@
| Knowledge Editing | Finetuning, LoRA [34], MEMIT [20] | zsRE [35] | Edit Success, Edit Generalization, Strict Edit Locality, MMLU | Order-free Error (Equation 1) Order Sensitivity (Equation 2) |
@@
| Model Compression | Pruning: Wanda [36], SparseGPT [16] Quantization: GPTQ [15], AWQ [37] | – | MMLU | |
@@
| Machine Unlearning | Gradient Ascent (GA) [38], Gradient Difference (GD) [39], Representation Misdirection Unlearning (RMU) [40] | WMDP [40] | WMDP, MMLU |
@@
Knowledge Editing. LMs’ knowledge can be fundamentally incorrect or grow outdated. Knowledge editing aims to update facts according to LMs, typically framed as question–answering. Here, where models are edited to return updated answers to questions by increasing the likelihood of the correct answers. We use three editors, including the popular MEMIT [20], as well as LoRA [34] and finetuning, which have recently been shown to be surprisingly successful editors despite their simplicity [41, 42, 43]. We use the zsRE [35] dataset, which is a popular question-answering benchmark for knowledge editing.
@@
To evaluate editors’ performance, we use three standard metrics [44]: Edit success measures whether the post-edit model successfully outputs the correct, edited response. Edit generalization measures how well the edit generalizes to other semantically-equivalent inputs. Following recent work, we compute generalization as the average edit success across a set of 10 holdout edit rephrasings. Strict Edit locality measures the impact of edits on unrelated inputs as the average success on random, unedited samples. All editing metrics are computed as the F1 score on the correct output logits as a strict measure of editing performance—intended outputs must be the most-likely tokens after editing. Each metric’s highest possible value is 1.0.
@@
Machine Unlearning. LMs can learn undesirable knowledge from their pretraining data, including information that is potentially dangerous [40] copyrighted [45, 46], toxic [47, 48, 9], or memorized [49, 50, 51, 52]. Machine unlearning methods aim to remove undesirable knowledge from LMs without regressing performance on extraneous domains [53, 27]. We experiment with three popular and representative methods: Gradient Ascent (GA) [38], Gradient Difference (GD) [39], and Representation Misdirection for Unlearning (RMU) [40]. GA is a traditional approach that minimizes the likelihood of chosen unlearning targets. GD augments GA by also maintaining the LM’s predictions on unrelated retention data. RMU, the most recent method, perturbs the model’s representations for inputs related to the learning target. Further details are in Appendix C.4.
@@
We evaluate unlearning with the recent QA dataset Weapons of Mass Description Proxy (WMDP) [40], updating LMs to forget potentially hazardous biosecurity and cybersecurity knowledge. We average the performance on WMDP’s cyber and bio splits, totaling 3,260 questions. We report individual split performance in Appendix D. Optimal unlearning yields 25% accuracy as this is random performance on the four-choice benchmark.
@@
Model Compression. LMs require significant compute resources during pretraining, finetuning, and inference. Model compression methods aim to reduce resources needed at test-time. We experiment with four popular, state-of-the-art weight pruning and quantization methods. For pruning, we choose SparseGPT [16] and Wanda [36]. For quantization, we choose GPTQ [15] and AWQ [37]. Pruning and quantization are used for different reasons, so we generally avoid comparisons between the two approaches. Each algorithm has a hyperparameter that controls how much to compress the model—% of zeroed weights for pruning, bits for quantization. Naturally, efficiency grows alongside compression, but at the expense of task performance. We vary this hyperparameter to explore the impacts of various levels of compression, choosing sparsity levels of .25, .35, .45, .55, .65, and .75 for pruning and 2, 3, 4, and 8 bits for quantization. The uncompressed model is 16 bits. Editing or unlearning some compressed models requires decompression, in which case we recompress afterwards using the same compression technique.
@@
Measuring General Model Utility. All interventions aim to avoid degrading overall model utility. While assessing general performance of LMs is an active area [54, 55, 56], we make the standard choice to evaluate question–answering accuracy on MMLU [57] using the LM Eval Harness [58]. Accuracy of 25% indicates random predictions.
@@ 4.1 Composing Model Compression with Knowledge Editing
Model compression degrades editing performance. Across Figure 2, we see that edits made by MEMIT, LoRA, and Finetuning deteriorate as compression increases. In weight pruning, the editing metrics, Success, Generalization, and Locality, steadily drop with higher sparsity. For quantization, the steepest drops occur between 2 and 4 Bits, and GPTQ decays editing performance faster than AWQ. While general performance decay from compression is unsurprising, the steepness of the decay differs dramatically between editors and compression methods. Decay curves also depend on the order of interventions. We also note that when compression is applied after editing, the only thing that can impact the editing metrics is the compression. Interestingly, Finetuning is also a surprisingly strong baseline at higher levels of compression, often achieving better editing performance than MEMIT or LoRA. Overall, these results suggests that model compression can limit a model’s editability.
@@
Editing performance hinges on the order of interventions. The shaded area between the curves in Figure 2 shows the impact of ordering. Interestingly, the ordering is highly impactful in most cases and varies dramatically between editing methods. For pruning methods (left side), compressing first is generally better than editing first. For example, pruning after Finetuning or LoRA destroys their edits much faster than if the model had been compressed first. For quantization (right side), the order also clearly matters. However, the best order differs by quantization method. For GPTQ, editing should be done first. For AWQ, compression should be done first. This finding suggests that these interventions interact in meaningful ways and successful composition will benefit from tailored, composable methods that have expected behaviors.
@@
Composability can vary within the same intervention category. We next summarize the general composability of each pair of interventions in Table 4.1, using the metrics proposed in Section 2: Order-free Error () measures best-case composition error, while Order Sensitivity () measures unwanted effects of intervention ordering. These metrics let us compare methods according to composability, and we find that 1) MEMIT is generally the most composable editor, totaling 14 wins compared to 9 by Finetuning and 4 by LoRA, 2) SparseGPT is the most composable pruning method with 8 wins compared to Wanda’s 6, and 3) AWQ far surpasses GPTQ, achieving 17 wins compared to GPTQ’s 2. However, many wins are quite close and composability can change dramatically between metrics. For example, Finetuning has better Order-free Error than MEMIT on 2/3 metrics but underforms MEMIT on Order Sensitivity for all metrics, so MEMIT’s edits are more robust. These results indicate a need to use many metrics to explicitly target and evaluate composability.
@@
Overall utility evaluations fail to measure composability. Despite similar MMLU scores, methods vary significantly in composability. For example, Table 4.1 shows that MEMIT and Wanda achieve near perfect Order Sensitivity on MMLU but have 0.24 Order Sensitivity on Edit Success. This is further backed by Figure 2, where all compositions look very similar according to MMLU, but not according to editing metrics. The disparity between overall utility and composability metrics underscores the need for more nuanced evaluations explicitly targeting composability as a design criterion.

### v2

@@ Abstract
Test-time interventions for language models can enhance factual accuracy, mitigate harmful outputs, and improve model efficiency without costly retraining.
But despite a flood of new methods, different types of interventions are largely developing independently.
In practice, multiple interventions must be applied sequentially to the same model, yet we lack standardized ways to study how interventions interact.
We fill this gap by introducing composable interventions, a framework to study the effects of using multiple interventions on the same language models, featuring new metrics and a unified codebase.
Using our framework, we conduct extensive experiments and compose popular methods from three emerging intervention categories—knowledge editing, model compression, and machine unlearning.
Our results over 417 different compositions uncover meaningful interactions: compression hinders editing and unlearning, composing interventions hinges on their order of application, and popular general-purpose metrics are inadequate for assessing composability.
Taken together, our findings showcase clear gaps in composability, suggesting a need for new multi-objective interventions.111All of our code is available at: github.com/hartvigsen-group/composable-interventions
*Equal contribution.
@@ 1 Introduction
Language models (LMs) exhibit striking capabilities on important tasks in many domains including medicine (Singhal et al., 2023), finance (Wu et al., 2023), science (Taylor et al., 2022), and entertainment (Zhong et al., 2023a). But despite high performance, deployed LMs can still misbehave unpredictably and require updates. For example, LMs generate content that is hallucinatory (Ji et al., 2023), factually incorrect (Zhao et al., 2023b), and harmful (Mendelsohn et al., 2023; Jain et al., 2024; Hartvigsen et al., 2022). Beyond unwanted behaviors, user requirements also change over time. For example, regulations arise (The President of the United States, 2023), computational resources constrict, knowledge gets outdated (Tack et al., 2024), and copyrighted training materials are identified (Grynbaum & Mac, 2023). Without ways to quickly address these issues, models can be left unsafe, outdated, biased, and non-compliant with laws or regulations, limiting their widespread responsible use (Kaddour et al., 2023).
@@
Many recent works study efficient, in-place updates for LMs. We broadly refer to these as interventions—updates to targeted properties of LMs applied after pretraining (and optional fine-tuning). For example, we can view model compression (Zhu et al., 2023; Frantar et al., 2023; Frantar & Alistarh, 2023) as an intervention to make language models more inference- or memory-efficient. Other rapidly-advancing examples include knowledge editing (Mazzia et al., 2023; Hartvigsen et al., 2023; Meng et al., 2022; 2023; Yu et al., 2024; Tan et al., 2024), detoxification (Welbl et al., 2021; Yu et al., 2023; Wang et al., 2022; Suau et al., ), and unlearning (Liu et al., 2024a; Lynch et al., 2024; Eldan & Russinovich, 2023; Lucki et al., 2024). However, these interventions are largely advancing independently. In practice, we usually have multiple requirements for properties of our models (e.g., factuality and efficiency). And as requirements change over time, new interventions must be applied. While some works have started studying interactions between training objectives (Matzken et al., 2023; Xu et al., 2023b; Li et al., 2024b), practical widespread use is limited without unified evaluations for how interventions interact.
@@
As depicted in Figure 1, we propose that practical interventions should be composable: When an intervention is applied to a model, it should avoid interfering with success of prior or future interventions. For example, a model user might need to repair a factual error in an LM using a knowledge editing method. Later on, they might have to quantize that model for deployment on mobile devices. A composable quantization method should specifically preserve the edited fact and a composable editor should withstand quantization. However, it remains unknown how well existing interventions compose, and we lack formal notions of intervention composition. We therefore propose two metrics for composability: 1) Order-free Error, where an intervention is composable if its application leaves others’ success unimpacted, and 2) Order Sensitivity, where the combined success of multiple interventions should not depend on the order in which they are applied. We instantiate these metrics in a codebase that includes popular interventions, tackling a major lack of standardized code that has inhibited cross-intervention evaluations until now.
@@
We use our framework to extensively study composability of state-of-the-art knowledge editing, model compression, and machine unlearning interventions on popular and recent LMs. Our experiments with 417 different compositions unearth novel insights and offer guidance for composing recent methods. Three key results are as follows: Model compression often limits the success of other interventions. This suggests a significant drawback of existing compression methods, which are nearly universal in LM deployments. The order in which interventions are applied dramatically alters their success. Therefore, we need new interventions explicitly designed for composability. General-purpose post-intervention model performance is a poor proxy for composability. This indicates that targeting composability as a metric has the potential to drive the development of new, practically-grounded interventions. Taken alongside our other findings (Section 4), our experiments suggest we sorely need to broaden intervention method evaluations and to design new methods explicitly for composability.
@@
We introduce the notion of composability for post-training interventions to LMs. Our work opens doors to address crucial, practical challenges in online LM updates and broadens evaluation criteria for interventions.
@@
2 Methods to Evaluate Intervention Composability
@@ 2.1 Preliminaries: Interventions for Language Models
Assume we are given a language model , where contains its pre-trained parameters and is the parameter space. We define an intervention as the combination of four things: 1) an operator that produces a new model , 2) hyperparameters , which typically control the strength of intervention, 3) a loss that specifies the targeted update,222Some interventions require data, but we remove data-dependent notation for simplicity. and 4) evaluation criteria , where higher scores indicate better performance. Each intervention’s operator aims to minimize loss and ultimately achieve high performance according to . For readability, we denote the application of intervention to model as .
@@ 2.2 Composable Interventions for Language Models
We propose that multiple interventions should compose with one another. Consider possible interventions . Intervention composition is the application of multiple interventions to a model in sequence.333For readability, we describe our framework in terms of two interventions, even though it scales. For example, we might apply intervention to model , then to the intervened-upon model. We can write this composition as either or .
Most interventions are developed specifically for their own criterion . For simplicity, we let each intervention have one criterion, though in practice many have multiple. We propose that the quality of many interventions in practice should also relate to how an intervention impacts our ability to apply other interventions. To meet this need, we propose two metrics that measure such composability. Each is computed for a given criterion and can be computed for a single hyperparameter setting or ranged over multiple settings to attain a richer description of intervention interactions.
@@
A smaller area again indicates better composability. And again, if hyperparameters are already chosen, this simplifies to a straightforward error: .
@@ 3.1 Intervention Datasets, Methods, and Metrics
We select ten methods from the three categories of interventions below. Each features task-specific and overall utility metrics. Appendix B further elaborates and Table 1 summarizes our setup.
| Interventions | Methods | Datasets | Intervention Metrics | Composability Metrics | Models |
@@
| Knowledge Editing | Finetuning, LoRA , MEMIT | zsRE | Edit Success, Edit Generalization, Strict Edit Locality, MMLU | Order-free Error (Equation 1) Order Sensitivity (Equation 2) | Llama-3 8B Mistral 7B Instruct Yi 1.5 9B Chat |
@@
| Model Compression | Pruning: Wanda, SparseGPT Quantization: GPTQ, AWQ | – | MMLU | ||
@@
| Machine Unlearning | Gradient Ascent (GA) , Gradient Difference (GD), Representation Misdirection Unlearning (RMU) | WMDP | WMDP, MMLU |
@@
Knowledge Editing. LMs’ knowledge can be fundamentally incorrect or grow outdated. Knowledge editing aims to update facts according to LMs, typically framed as question–answering. Here, where models are edited to return updated answers to questions by increasing the likelihood of the correct answers. We use three methods, including the popular MEMIT (Meng et al., 2023) editor, as well as LoRA (Hu et al., 2021) and finetuning, which have recently been shown to be surprisingly successful editors, despite their simplicity (Gangadhar & Stratos, 2024; Hsueh et al., 2024; Hua et al., 2024). We use the zsRE (Levy et al., 2017) dataset, which is a popular question-answering benchmark for knowledge editing.
@@
To evaluate editors’ performance, we use three standard metrics (Yao et al., 2023): Edit success measures whether the post-edit model successfully outputs the correct, edited response. Edit generalization measures how well the edit generalizes to other semantically-equivalent inputs. We compute generalization as the average edit success across a set of 10 holdout edit rephrasings. Strict Edit locality measures the impact of edits on unrelated inputs as the average success on random, unedited samples. All editing metrics are computed as the F1 score on the correct output logits as a strict measure of editing performance—intended outputs must be the most-likely tokens after editing. Each metric’s highest value is 1.0.
@@
Machine Unlearning. LMs can learn undesirable knowledge from their pretraining data, including information that is potentially dangerous (Li et al., 2024a) copyrighted (Karamolegkou et al., 2023; Meeus et al., 2024), toxic (Sheng et al., 2019; Gehman et al., 2020; Hartvigsen et al., 2022), or memorized (Carlini et al., 2018; Ippolito et al., 2022; Biderman et al., 2023; Prashanth et al., 2024). Machine unlearning methods aim to remove undesirable knowledge from LMs without regressing performance on extraneous domains (Cao & Yang, 2015; Liu et al., 2024a). We experiment with three popular and representative methods: Gradient Ascent (GA) (Jang et al., 2022), Gradient Difference (GD) (Maini et al., 2024), and Representation Misdirection for Unlearning (RMU) (Li et al., 2024a). GA is a traditional approach that minimizes the likelihood of chosen unlearning targets. GD augments GA by also maintaining the LM’s predictions on unrelated retention data. RMU, the most recent method, perturbs the model’s representations for inputs related to the learning target. Further details are in Appendix C.4.
@@
We evaluate unlearning with Weapons of Mass Description Proxy (WMDP) (Li et al., 2024a), updating LMs to forget potentially hazardous biosecurity and cybersecurity knowledge. We average the performance on WMDP’s cyber and bio splits, totaling 3,260 questions. We report individual split performance in Appendix E. Optimal unlearning yields 25% accuracy as this is random performance on the four-choice benchmark.
Unlearning is a broad problem area. While, WMDP is among the most studied unlearning benchmarks, recent works have suggested that WMDP performance provides an imperfect evaluation of unlearning writ-large. We proceed with leveraging WMDP as it is a sufficient evaluation setting for measuring composability when applied to unlearning, but caveat that comprehensive evaluation of unlearning is an open research problem.
@@
Model Compression. LMs require significant compute resources during pretraining, finetuning, and inference. Model compression methods aim to reduce resources needed at test-time. We experiment with four popular, state-of-the-art weight pruning and quantization methods. For pruning, we choose SparseGPT (Frantar & Alistarh, 2023) and Wanda (Sun et al., 2023). For quantization, we choose GPTQ (Frantar et al., 2023) and AWQ (Lin et al., 2023). Pruning and quantization are used for different reasons, so we generally avoid comparisons between the two. Each algorithm has a hyperparameter that controls the compression level—% of zeroed weights for pruning, bits for quantization. Naturally, efficiency grows with compression, but at the expense of task performance. We vary this hyperparameter to explore various levels of compression, choosing sparsity levels of .25, .35, .45, .55, .65, and .75 for pruning and 2, 3, 4, and 8 bits for quantization – down from 16 bits. We note that most compression techniques require decompressing models, so we recompress after editing or unlearning using the same compression technique.
@@
Measuring General Model Utility. All interventions aim to avoid degrading overall model utility. While assessing general performance of LMs is an active area (Dehghani et al., 2021; Biderman et al., 2024; Gupta et al., 2024), we make the standard choice to evaluate question–answering accuracy on MMLU (Hendrycks et al., 2020) using the LM Eval Harness (Gao et al., 2023). Accuracy of 25% indicates random predictions.
@@ 3.2 Implementation details
All experiments in our main results (Section 4) are performed with Llama3-8B (AI@Meta, 2024), which is well-studied and highly capable for its parameter count. All results for knowledge editing methods are averaged over 10 batches of 50 randomly-selected edits from zsRE. Further details on implementations and hyperparameters are in Appendix C.
@@ 4.1 Composing Model Compression with Knowledge Editing
Model compression degrades editing performance. Across Figure 2, edits made by MEMIT, LoRA, and Finetuning deteriorate as compression increases. In weight pruning, the editing metrics, Success, Generalization, and Locality, steadily drop with higher sparsity. For quantization, the steepest drops occur between 2 and 4 Bits, and GPTQ decays editing performance faster than AWQ. While general performance decay from compression is unsurprising, its steepness varies dramatically between editors and compression methods. Decay curves also depend on the order of interventions. Finetuning remains a surprisingly strong baseline at higher compression levels, often surpassing MEMIT or LoRA. Overall, these results suggests that model compression can limit a model’s editability.
@@
Editing performance hinges on the order of interventions. The shaded area between the curves in Figure 2 shows the impact of ordering. Ordering is highly impactful in most cases and varies dramatically between editing methods. For pruning methods (left side), compressing first is generally better than editing first. For example, pruning after Finetuning or LoRA destroys edits much faster than if the model had been compressed first. For quantization (right side), the order also matters, but the best order differs by method. For GPTQ, editing should be done first. For AWQ, compression should be done first. This finding suggests that these interventions interact in meaningful ways and successful composition will benefit from tailored, composable methods that have expected behaviors.
@@
Composability can vary within the same intervention category. We next summarize the general composability of each intervention pair in Table 4.1, using the metrics from Section 2: Order-free Error () measures best-case composition error, while Order Sensitivity () measures unwanted effects of intervention ordering. These metrics let us compare methods according to composability, and we find that 1) MEMIT is generally the most composable editor, totaling 14 wins compared to 9 by Finetuning and 4 by LoRA, 2) SparseGPT is the most composable pruning method with 8 wins compared to Wanda’s 6, and 3) AWQ far surpasses GPTQ, achieving 17 wins compared to GPTQ’s 2. However, many wins are quite close and composability can change dramatically between metrics. These results indicate a need to use many metrics to explicitly target and evaluate composability.
@@
Overall utility evaluations fail to measure composability. Despite similar MMLU scores, methods vary significantly in composability. For example, Table 4.1 shows that MEMIT and Wanda achieve near-perfect Order Sensitivity on MMLU but have 0.24 Order Sensitivity on Edit Success. This is bolstered by Figure 2, where all compositions are similar according to MMLU, but not according to editing metrics. The disparity between overall utility and composability metrics underscores the need for more nuanced evaluations explicitly targeting composability as a design criterion.